# Evidence for simple “arrow of time functions” in closed chaotic quantum systems

### v1

@@ Abstract
Through an explicit construction, we assign to any infinite temperature autocorrelation function $C(t)$ a set of functions $\alpha^{n}(t)$. The construction of $\alpha^{n}(t)$ from $C(t)$ requires the first $2n$ temporal derivatives of $C(t)$ at times 00 and $t$. Our focus is on $\alpha^{n}(t)$ that (almost) monotonously decrease, we call these “arrows of time functions” (AOTFs). For autocorrelation functions of few body observables we numerically observe the following: An AOTF featuring a low $n$ may always be found unless the the system is in or close to a nonchaotic regime with respect to a variation of some system parameter. All $\alpha^{n}(t)$ put upper bounds to the respective autocorrelation functions, i.e. $\alpha^{n}(t)\geq C^{2}(t)$. Thus the implication of the existence of an AOTF is comparable to that of the H-Theorem, as it indicates a directed approach to equilibrium. We furthermore argue that our numerical finding may to some extent be traced back to the operator growth hypothesis. This argument is laid out in the framework of the so-called recursion method.
@@ I Introduction
Is there a tendency in closed, nondriven quantum systems to approach equilibrium in a monotonous fashion? And if so, how would one even pinpoint all the somewhat vague notions in this question? Obviously this question is related to the question for the so called “arrow of time”. A full fledged overview over all approaches to this subject, even restricted to the quantum realm, is clearly beyond the possibilities of the paper at hand. Thus, below we only very briefly introduce some concepts that appear to be in the vicinity/form the background of what shall be mainly presented later in the present paper. Readers well acquainted with these subjects may just skip this part.
@@
arrow of time, observational entropy and consistent histories: This approach may be regarded as an synthesis of the open and the closed system concepts. It is based on the observation that some dynamics of a (typically more complex) closed system may appear as if the latter was subject to external measurements. In this case so called “consistent histories” emerge. Whether or not this is the case could in principle be answered by computing a so called “decoherence functional” [10], in practice this may be hard. However, if consistent histories emerge, the results from the open quantum quantum system carry over and a suitably defined analog of the von Neumann entropy is nonincreasing [11]. It has also been pointed out that in this case an even more accessible quantity called “observational entropy” [12, 13, 14] is nonincreasing. While these are promising perspectives, it is unclear whether consistent histories always emerge when the notion of an arrow of time appears applicable.
In the paper at hand we are working within the closed system framework. However, we aim at going beyond the common lore of “scarceness of excursions from equilibrium” towards establishing a time-local arrow of time in the spirit of the Boltzmann H-theorem [15]. While a substantial part of this effort consists in numerical observations, we attempt to trace these finding partially back to the operator growth hypothesis [16] in the second part of the present paper. Throughout this analysis we take the “chaoticity” of a system into account which we measure by the gap ratio [17].
@@ II Construction of arrow of time functions
$\begin{split}\tilde{\varphi}_{j}(t)&:=\partial_{t}\varphi_{j-1}(t)-b_{j-1}\varphi_{j-2}(t),\\ b_{j}^{2}&:=\langle\tilde{\varphi}_{j}(t),\tilde{\varphi}_{j}(t)\rangle,\\ \varphi_{j}(t)&:=-b_{j}^{-1}\tilde{\varphi}_{j}(t),\end{split}$ (1)
with $b_{0}=0$. The inner product $\langle$-,$\rangle$ - , - ⟩ is defined on the basis $\{e_{k}\}=\{\partial_{t}^{k}C(t)\}$ via $\langle e_{n},e_{m}\rangle:=|\partial_{t}^{n+m}C(t)|_{t=0}|$ and constitutes well-defined scalar product on the space spanned by the first $M$ derivatives of $C(t)$. With these $\{\varphi_{j}(t)\}$ we construct
@@
We then determine $\mathcal{I}$ as the smallest value for which $\alpha^{\mathcal{I}}$ becomes monotonously decreasing up to a total ascend $\delta$ until the relaxation time $\tau_{\mathcal{I}}$ , the latter being defined by $\alpha^{\mathcal{I}}(t)\leq\delta\quad\forall t\geq\tau_{\mathcal{I}}$. (For the issue of Poincaré recurrences, cf. Ref.  [^1] Strictly speaking such a time $\tau_{\mathcal{I}}$ does not exist for any finite system, since there will always be Poincaré recurrences. These, however occur by magnitudes later than any time scale considered in the paper at hand. Technically $\tau_{\mathcal{I}}$ is determined from $\alpha^{\mathcal{I}}(t)\leq\delta\quad\forall t\geq\tau_{\mathcal{I}}\geq T$ where $T$ is the largest time that we can numerically access. Thus $T$ varies with the considered examples but is always substantially larger than $\tau_{\mathcal{I}}$.) We also refer to the quantity $\mathcal{I}$ as the intricacy of $\alpha$. The value of $\delta$ should be suitably chosen in a sense that is supposed to become clear in the remainder of this paper. For the entirety of the paper at hand we set $\delta=1/200$. For given $C(t),\delta$ a finite $\mathcal{I}$ may or may not exist. We interpret $\alpha^{\mathcal{I}}(t)$ as a deviation-from-equilibrium, e.g., since it limits possible values of $C(t)$ through $\alpha^{\mathcal{I}}(t)\geq C^{2}(t)$. Due to its almost monotonous decrease, we also refer to $\alpha^{\mathcal{I}}(t)$ as an arrow of time function (AOTF). The degree to which the function $\alpha^{\mathcal{I}}(t)$ may be described as time-local depends on $\mathcal{I}$. From the construction one readily infers that $2\mathcal{I}$ temporal derivatives at $t=0$ and $\mathcal{I}$ temporal derivatives at $t$ are required to actually compute $\alpha^{\mathcal{I}}(t)$. Thus, for not too large $\mathcal{I}$, we refer to $\alpha^{\mathcal{I}}(t)$ as time-local with respect to $C(t)$. Physically numerous but in this context somewhat trivial cases emerge for autocorrelation functions $C(t)$ which are themselves monotonously decaying, regardless of the decay time. In this case $\alpha^{0}(t)$ is an AOTF. While this is by no means excluded, this investigation focuses on more complex $C(t)$ that do not decay monotonously but induce AOTFs with finite $\mathcal{I}$. For first examples of AOTFs, see Figs. 1. Note that $C(t)$ in Fig. 1 (bottom) is, unlike all other examples in this paper, not specifically computed from a many-body quantum model, but simply conjectured. Nevertheless a corresponding AOTF may be constructed according to Eqs. (1,2).
@@
To contextualize AOTFs we briefly comment on their comparability to observational entropy [12, 13]. Let $C(t)$ correspond to the (infinite temperature) autocorrelation function of some observable $\mathcal{O}$. Let furthermore $\rho\propto\exp(\beta\mathcal{O})$ with small $\beta$. Then $\alpha^{\mathcal{I}}(t)$ may be computed from the expectation values $\langle\partial_{t}^{n}\mathcal{O}(t)\rangle$ with $0\leq n\leq\mathcal{I}$. Thus, just like observational entropy is based on the knowledge of a specific, case dependent set of expectation values, so is $\alpha^{\mathcal{I}}(t)$. Note however that the $\langle\partial_{t}^{n}\mathcal{O}(t)\rangle$ must not be interpreted as probabilities, as neither $\langle\partial_{t}^{n}\mathcal{O}(t)\rangle>0$ nor $\sum_{n}\langle\partial_{t}^{n}\mathcal{O}(t)\rangle=1$ holds. Thus, the concept of the AOTFs presented here should be considered as complementary to the concept of observational entropy.
Figure 1: Top: Sketch of the functions $\alpha^{n}(t)$ as outlined in (2) for different values for $n$ for the $\sigma_{S}^{x}$ observable in the spin-bath model (16) with $\lambda=0.7$. The index $n$ increases until the construction becomes decreasing up to $\delta$ for some $n=:\mathcal{I}$ (here: $\mathcal{I}=2$). Each $\alpha^{n}(t)$ upper-bounds its predecessor $\alpha^{n-1}(t)$. Inset: The respective autocorrelation function $C(t)$. Bottom: AOTF for the example of robust revivals put forward in Ref. [4]. Concretely we have $C(t)=e^{-5t^{2}}+0.5\left(e^{-5(t-3.5)^{2}}+e^{-5(t+3.5)^{2}}\right)$, see inset. Regardless of the apparent equilibration prior to the revival around $t_{\text{rev}}\sim 3.5$ the AOTF correctly identifies the systems’ eventual approach to equilibrium by giving rise to a monotonous function. Here we find $\mathcal{I}=28$.
III Emergence of arrows of time
@@
where $s_{k}=\epsilon_{k+1}-\epsilon_{k}$ with the $\epsilon_{k}$ as the eigenenergies of the considered system. Exploiting respective symmetries, we truncate the system to a specific symmetry separated subspace. For fully chaotic systems, the mean gap ratio is numerically determined as $\langle r\rangle_{\text{GOE}}\approx 0.53$ by means of large random matrices from a Gaussian orthogonal ensemble (GOE). Conversely, for integrable systems exhibiting a Poissonian spectrum, $\langle r\rangle_{P}\approx 0.386$ [17].
Whilst tracking the respective gap ratio along $\sigma$ we also compute the autocorrelation $C(t)$ for different system parameters. For these autocorrelation functions we determine $\mathcal{I}$ as the intricacy of the arrow of time function as described in Sec. II.
For details on the numerical method applied in this work we refer to Appendix A.
@@ III.1 Models and observables
Two-chain model: As a first model we consider two (mixed-field) Ising chains with local coupling of the respective $\sigma^{z}$-terms moderated by the interaction strength $\lambda$ and with periodic boundary conditions. The Hamiltonian reads
@@
As the model mimics the pertinent scenario of the equilibration of two macroscopic bodies in contact, we consequently consider the energy imbalance
@@
For low $\lambda$ an (almost) exponentially decaying $C(t)$ results, see Fig. 2. This is reminiscent of the decay of a temperature difference between to macroscopic bodies in mutual thermal contact, but insulated otherwise. The application of a standard projection operator technique in the Markovian, weak coupling limit would find this result [7]. The intricacy is $\mathcal{I}=0$. However, as we increase $\lambda$ we leave the Markovian regime and at $\lambda\sim 2$ $C(t)$ is no longer monotonously decaying but exhibits a more intricate form, see Fig. 2. Nevertheless the intricacy is moderate , $\mathcal{I}=1$ and $\langle r\rangle$ is still close the GOE value, cf. Fig. 3. Only if $\lambda$ is increased even further towards $\lambda\sim 2.5$ the gap ratio $\langle r\rangle$ sinks below the GOE value (since the model approaches a situation of decoupled dimers) and the intricacy substantially increases to $\mathcal{I}=86$.
Figure 2: Autocorrelation functions $C(t)$ (left) and their respective squares $C^{2}(t)$ and AOTFs $\alpha^{\mathcal{I}}$ (right) for energy imbalance in the two-chain-model (4) for different interaction strengths $\lambda$. Top: At $\lambda=0.7$ (chaotic regime, see Fig. 3) $C(t)$ itself is monotonously decreasing and hence  $\alpha^{0}$ is an AOTF . Centre: At $\lambda=2.2$ (chaotic regime, see Fig. 3) $C(t)$ no longer decreases monotonously. However, $\alpha^{1}$ does, thus two derivatives w.r.t. time are required to construct an AOTF . Bottom: In the nonchaotic regime at $\lambda=2.7$ (see Fig. 3) a total of 260 derivatives w.r.t. time are required to construct the AOTF $\alpha^{130}$.
In the direction of even larger $\lambda$, i.e. leaving the fully chaotic regime, we observe further growth of $\mathcal{I}$, see Fig. 3. For this model we can rule out arrows of time for $\mathcal{I}\leq 200$ for $\lambda\gtrsim 2.8$. Conversely, in the chaotic regime we find arrows of time with low intricacy giving rise to a time-local description of the approach to equilibrium as only a low finite numbers of derivatives of the autocorrelation function $C(t)$ at times 00 and $t$ are sufficient to construct an AOTF.
@@
in which we set $(J,\Delta^{\prime})=(1,0.5)$ and where $s_{k}^{a}=\sigma_{k}^{a}/2$ denote the spin-matrices. For the entirety of this paper we set $\hbar=1$. We vary the nearest-neighbour interaction term $\Delta$. Introducing the next-nearest interaction $\Delta^{\prime}$ breaks Bethe-Ansatz integrability [23, 24]. However, for $\Delta\gg\Delta^{\prime}$ the system again approaches the integrable regime, see Fig. 5. For any isotropies $\Delta,\Delta^{\prime}$ the total magnetization is conserved [23], motivating the study of the respective current
@@
Figure 4: Autocorrelation functions $C(t)$ (left) and their respective squares $C^{2}(t)$ and AOTFs $\alpha^{\mathcal{I}}$ (right) for magnetisation current in the XXZ model (9) for different nearest-neighbour interactions $\Delta$. Top: At $\Delta=3$ (chaotic regime, see Fig. 5) $C(t)$ does not decay monotonously but the AOTF $\alpha^{3}(t)$ does. Bottom: Further away from the chaotic regime, at $\Delta=10$, an AOTF is reached only at $\mathcal{I}=47$, i.e. $\alpha^{47}(t)$ decays monotonously.
@@
Lastly, we include to our investigation the energy current
@@
where we choose the parameters as $(\omega,J,B_{x},B_{z})=(1,1,1,0.5)$. Varying the coupling strength $\lambda$ we compute the autocorrelation for the $x$ and $z-$component of coupled spin respectively, i.e. $\mathcal{O}^{(x)}=\sigma_{S}^{x}$ and $\mathcal{O}^{(z)}=\sigma_{S}^{z}$. See Figs. 1 and 8 for exemplary AOTFs.
Figure 8: Autocorrelation functions $C(t)$ (left) and their respective squares $C^{2}(t)$ and AOTFs $\alpha^{\mathcal{I}}$ (right) for the observable $\mathcal{O}_{S}^{z}$ in the spin bath model (16) for different couplings $\lambda$. Top: At small coupling $\lambda$, here: $\lambda=2$ (chaotic regime, see Fig. 9) we find an AOTF with intricacy $\mathcal{I}=10$. Bottom: Approaching the chaotic regime again at $\lambda=12$, the intricacy of the respective AOTF drops to $\mathcal{I}=48$ again falling within our numerical scope.
@@
At small coupling $\lambda$ the intricacies $\mathcal{I}$ for the AOTFs of $\sigma_{S}^{z}$ start off at $\mathcal{I}=0$ as at first the autocorrelation functions are monotonous themselves. With increasing coupling strength, the intricacies grow until $\mathcal{I}=19$ for $\lambda=2.5$. As the system approaches the nonchaotic regime the intricacies exceed $\mathcal{I}=100$, leaving the scope of our numerical method. Once the system approaches the chaotic regime again at $\lambda\sim 11$ we again infer intricacies $\mathcal{I}<100$ starting at $\lambda=12$ where $\mathcal{I}=48$. These then decrease further as the system moves away from the nonchaotic regime, eventually giving rise to an AOTF with $\mathcal{I}=5$ at $\lambda=20$.
In the case of the observable $\sigma_{S}^{x}$ the picture is different. The intricacies do not seem to be affected by the systems’ transition from the chaotic to an nonchaotic regime. With an exception at $\lambda=2$ where our construction finds the respective autocorrelation function to be suitably monotonous, i.e. $\mathcal{I}=0$, we find AOTFs with an intricacy $\mathcal{I}=2$ for any other considered coupling strength $\lambda\leq 8.5$. Entering a strong coupling range at $\lambda\gtrsim 9$ however, the functions $C^{2}(t)$ do no longer decay within a reasonable period in the sense we require to infer an equilibration time $\tau$. Consequently we leave them out of further discussions.
Having presented all numerical evidence, we now aim at distilling the common principle that is present in all these examples: Whenever an autocorrelation function of a few body observable exhibits high intricacy, e.g. $\mathcal{I}>20$ (for the $\delta$ chosen here) the corresponding system either belongs to the nonchaotic regime (according to the gap ratio) or a nonchaotic regime is close by with respect to variation of some model parameter. Note that the converse does not hold: low intricacy dynamics may emerge in nonchaotic regimes, see e.g. $\sigma_{S}^{z}$ in the spin-bath model. This principle is our main numerical result.

### v3

@@ Abstract
Through an explicit construction, we assign to any infinite temperature autocorrelation function $C(t)$ a set of functions $\alpha^{n}(t)$. The construction of $\alpha^{n}(t)$ from $C(t)$ requires the first $2n$ temporal derivatives of $C(t)$ at times 00 and $t$. Our focus is on $\alpha^{n}(t)$ that (almost) monotonously decrease; we call these “arrows of time functions” (AOTFs). For autocorrelation functions of few-body observables we numerically observe the following: An AOTF featuring a low $n$ may always be found unless the system is in or close to a nonchaotic regime with respect to a variation of some system parameter. All $\alpha^{n}(t)$ put upper bounds to the respective autocorrelation functions, i.e. $\alpha^{n}(t)\geq C^{2}(t)$. Thus the implication of the existence of an AOTF is comparable to that of the H-Theorem, as it indicates a directed approach to equilibrium. We furthermore argue that our numerical finding may to some extent be traced back to the operator growth hypothesis. This argument is laid out in the framework of the so-called recursion method.
@@ I Introduction
Is there a tendency in closed, nondriven quantum systems to approach equilibrium in a monotonous fashion? And if so, how would one even pinpoint all the somewhat vague notions in this question? Obviously this question is related to the so-called “arrow of time”. A full-fledged overview over all approaches to this subject, even restricted to the quantum realm, is clearly beyond the possibilities of the paper at hand. Thus, below we only very briefly introduce some concepts that appear to be in the vicinity or form the background of what shall be mainly presented later in the present paper. Readers well acquainted with these subjects may just skip this part.
@@
Arrow of time, observational entropy, Boltzmann entropy and consistent histories: This approach may be regarded as a synthesis of the open and the closed system concepts. It is based on the observation that some dynamics of a (typically more complex) closed system may appear as if the latter was subject to external measurements. In this case so-called “consistent histories” emerge. Whether or not this is the case could in principle be answered by computing a so-called “decoherence functional” [10], in practice this may be hard. However, if consistent histories emerge, the results from the open quantum system carry over and a suitably defined analog of the von Neumann entropy is nondecreasing [11]. It has also been pointed out that in this case an even more accessible quantity called “observational entropy” [12, 13, 14] is nondecreasing. It has also been shown that for an ideal quantum gas the Boltzmann entropy, a quantity conceptually similar to observational entropy, is nondecreasing [15]. These approaches are closest to the main subject of the present paper, as they target monotonous
(entropy-)functions that are entirely defined within closed, non-driven systems. However, while the “arrow of time functions” (which will be defined below) are no thermodynamic state functions, their monotonicity appears to prevail in “pure quantum settings” where no consistent histories emerge. The latter does not hold for observational entropy [14].
In the paper at hand we are working within the closed-system framework. However, we aim at going beyond the common lore of “scarceness of excursions from equilibrium” towards establishing a time-local arrow of time in the spirit of the Boltzmann H-theorem [16]. While a substantial part of this effort consists in numerical observations, we attempt to trace these finding partially back to the operator growth hypothesis [17] in the second part of the present paper. Throughout this analysis we take the “chaoticity” of a system into account which we measure by the gap ratio [18].
@@
(Some) other approaches to irreversibility in closed quantum systems: Several considerations suggest that autocorrelation functions in chaotic (quantum) systems are well expandable in terms of (a few) exponentials with negative real parts in the exponents, either from the beginning, or at least in the long run [23, 24, 25, 26]. This concept goes under the name of Ruelle-Pollicott resonances, pseudomodes or quasinormal modes. If such an approach applies a direction of time is naturally established.
Another line of research is based on the so-called volume or Hertz entropy. The latter features various thermodynamically desirable properties. There is also evidence that it is nondecreasing in certain scenarios [27, 28, 29, 30].
@@ II Construction of arrow of time functions
$\begin{split}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\varphi}_{j}(t)}&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:=-\partial_{t}\varphi_{j-1}(t)+b_{j-1}\varphi_{j-2}(t)},\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}b_{j}^{2}}&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:=\langle\tilde{\varphi}_{j}(t),\tilde{\varphi}_{j}(t)\rangle,}\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi_{j}(t)}&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:=b_{j}^{-1}\tilde{\varphi}_{j}(t),}\end{split}$ (1)
with $b_{0}=0$. The inner product $\langle$-,$\rangle$ - , - ⟩ is defined on the basis $\{\partial_{t}^{k}C(t)\}$ via $\langle\partial_{t}^{n}C(t),\partial_{t}^{m}C(t)\rangle:=(-1)^{n}\partial_{t}^{n+m}C(t)|_{t=0}$ and constitutes a well-defined scalar product on the space spanned by the first $M$ derivatives of $C(t)$. For further details, see App. A. With these $\{\varphi_{j}(t)\}$ we construct
@@
We then determine $\mathcal{I}$ as the smallest value for which $\alpha^{\mathcal{I}}$ becomes monotonously decreasing up to an accumulated error $\delta$ until the relaxation time $\tau_{\mathcal{I}}$  [^1] We infer the relaxation time $\tau_{\mathcal{I}}$ as the time after which the function $\alpha^{n}(t)$ has permanently fallen below some relaxation threshold. Concretely we determine $\tau_{\mathcal{I}}$ as the earliest time such that $\alpha^{n}(t)\leq 1/200$ $\forall\ t\geq\tau_{\mathcal{I}}$. The ”accumulated error” of a function $\alpha^{n}(t)$ therefore refers to the sum of all ascends of the function until $\tau_{\mathcal{I}}$. (For the issue of Poincaré recurrences, cf. Ref.  [^2] Strictly speaking such a time $\tau_{\mathcal{I}}$ does not exist for any finite system, since there will always be Poincaré recurrences. These, however occur by magnitudes later than any time scale considered in the paper at hand. Technically $\tau_{\mathcal{I}}$ is always substantially smaller than the largest time we can access numerically..) In order to practically determine $\mathcal{I}$ for some $C(t)$ we use numerical methods and simply check the monotonicity of $\alpha^{n}(t)$ while increasing $n$ “one by one”, see also Fig. 2 (lower-middle row). The value of $\delta$ should be suitably chosen in a sense that is supposed to become clear in the remainder of this paper. For the entirety of the paper at hand we set $\delta=1/200$. We refer to the quantity $\mathcal{I}$ as intricacy for reasons which will also become clear below. In passing we mention that the $\alpha^{n}(t)$ have an intrinsic connection to the “Krylov-complexity” [17] as detailed in Sect. IV.2.
@@
For given $C(t),\delta$ a finite $\mathcal{I}$ may or may not exist. We interpret $\alpha^{\mathcal{I}}(t)$ as a deviation-from-equilibrium, e.g., since it limits possible values of $C(t)$ through $\alpha^{\mathcal{I}}(t)\geq C^{2}(t)$. Due to its almost monotonous decrease, we also refer to $\alpha^{\mathcal{I}}(t)$ as an arrow of time function (AOTF). The degree to which the function $\alpha^{\mathcal{I}}(t)$ may be described as time-local depends on $\mathcal{I}$. From the construction one readily infers that $2\mathcal{I}$ temporal derivatives at $t=0$ and $\mathcal{I}$ temporal derivatives at $t$ are required to actually compute $\alpha^{\mathcal{I}}(t)$. Thus, for not too large $\mathcal{I}$, we refer to $\alpha^{\mathcal{I}}(t)$ as time-local with respect to $C(t)$. Physically numerous but in this context somewhat trivial cases emerge for autocorrelation functions $C(t)$ which are themselves monotonously decaying, regardless of the decay time. In this case $\alpha^{0}(t)$ is an AOTF. While this is by no means excluded, this investigation focuses on more complex $C(t)$ that do not decay monotonously but induce AOTFs with finite $\mathcal{I}$.
@@
To contextualize AOTFs we briefly comment on their comparability to observational entropy [12, 13] as both objects are closely related to irreversibility and defined within closed quantum systems. Let $C(t)$ correspond to the (infinite temperature) autocorrelation function of some observable $\mathcal{O}$. Consider an initial state $\rho\propto\exp(\beta\mathcal{O})$ with small $\beta$. Then $\alpha^{\mathcal{I}}(t)$ may be computed from the expectation values $\langle\partial_{t}^{n}\mathcal{O}(t)\rangle$ with $0\leq n\leq\mathcal{I}$. Thus, just like observational entropy is based on the knowledge of a case dependent set of expectation values, so is $\alpha^{\mathcal{I}}(t)$ in this case. Note, however, that the $\langle\partial_{t}^{n}\mathcal{O}(t)\rangle$ must not be interpreted as probabilities, as neither $\langle\partial_{t}^{n}\mathcal{O}(t)\rangle>0$ nor $\sum_{n}\langle\partial_{t}^{n}\mathcal{O}(t)\rangle=1$ holds. Furthermore, unlike (observational) entropies, the $\alpha^{\mathcal{I}}(t)$ are no thermodynamic state functions, nevertheless they indicate the direction of time in the dynamics of $C(t)$ (or, in the above special case $\langle\mathcal{O}(t)\rangle$) including the prediction of an eventual stationary value. Thus, the concept of the AOTFs presented here should be considered as complementary to the concept of observational entropy. Other than for observational entropy, the monotonicity of the $\alpha^{\mathcal{I}}(t)$ (with low $\mathcal{I}$) appears not to depend on slowness of the observable or the emergence of consistent histories, cf. Ref. [14, 33] as will become evident below.
III Numerical evidence for arrows of time
@@
where $s_{k}=\epsilon_{k+1}-\epsilon_{k}$ with the $\epsilon_{k}$ as the eigenenergies of the considered system. Exploiting respective symmetries, we truncate the system to a specific symmetry-separated subspace. For fully chaotic systems, the mean gap ratio is numerically determined as $\langle r\rangle_{\text{GOE}}\approx 0.53$ by means of large random matrices from a Gaussian orthogonal ensemble (GOE). Conversely, for integrable systems exhibiting a Poissonian spectrum, $\langle r\rangle_{P}\approx 0.386$ [18].
While tracking the respective gap ratio along $\sigma$, we also compute the autocorrelation $C(t)$ for different system parameters. For these autocorrelation functions we determine $\mathcal{I}$ as the intricacy of the arrow of time function as described in Sec. II.
For details on the numerical method applied in this work, we refer to Appendix B.
@@ III.1 Models and observables
Two-chain model: As a first model we consider two (mixed-field) Ising chains with local coupling of the respective $\sigma^{z}$-terms moderated by the interaction strength $\lambda$ and with periodic boundary conditions. For the entirety of this paper we set $\hbar=1$. The Hamiltonian reads
@@
Figure 1: Sketch of the archetypal example of two macroscopic bodies in contact (here: two cups of coffee) described by the Hamiltonian given by (4).
As the model mimics the pertinent scenario of the equilibration of two macroscopic bodies in contact (see Fig. 1), we consequently consider the energy imbalance
@@
For low $\lambda$ an (almost) exponentially decaying $C(t)$ results, see Fig. 2. This is reminiscent of the decay of a temperature difference between two macroscopic bodies in mutual thermal contact, but insulated otherwise. The application of a standard projection operator technique in the Markovian, weak-coupling limit would find this result [7]. The intricacy is $\mathcal{I}=0$. However, as we increase $\lambda$ we leave the Markovian regime and at $\lambda\sim 2$ $C(t)$ is no longer monotonously decaying but exhibits a more intricate form, see Fig. 2. Nevertheless the intricacy is moderate, $\mathcal{I}=1$ and $\langle r\rangle$ is still close to the GOE value, cf. Fig. 3. (For an assessment of the observational entropy approach to this scenario, see Ref.   [^3] We expect the emergence of consistent histories and Markovianity to be absent at (and above) $\lambda\approx 2$, hence we do not expect a pertinent observational entropy to be monotonously decreasing..) If $\lambda$ is increased further, the gap ratio $\langle r\rangle$ starts to sink below the GOE value (since the model approaches a situation of decoupled dimers) and the intricacy rapidly increases towards much higher values, see Figs. 2,3.
Figure 2: Autocorrelation functions $C(t)$ (left) and their respective squares $C^{2}(t)$ and AOTFs $\alpha^{\mathcal{I}}(t)$ (right) for the energy imbalance in the two-chain-model (4) for different interaction strengths $\lambda$. Top: At $\lambda=0.7$ (chaotic regime, see Fig. 3) $C(t)$ itself is monotonously decreasing and hence  $\alpha^{0}(t)$ is an AOTF . Upper-middle: At $\lambda=2.2$ (chaotic regime, see Fig. 3) $C(t)$ no longer decreases monotonously. However, $\alpha^{1}(t)$ does, thus two derivatives w.r.t. time are required to construct an AOTF. Lower-middle: At interaction strength $\lambda=2.4$ (nonchaotic regime, see Fig. 3) the intricacy $\mathcal{I}$ is 20. We also depict some intermediate functions $\alpha^{n}(t)$, (here only $n=5,10,15$ to avoid cluttering) that fall short of the required monotonicity standard explained in Sec. 1. Bottom: In the nonchaotic regime at $\lambda=2.7$ (see Fig. 3) a total of 260 derivatives w.r.t. time are required to construct the AOTF $\alpha^{130}(t)$.
For this model we can rule out arrows of time for $\mathcal{I}\leq 200$ for $\lambda\gtrsim 2.8$. Conversely, in the chaotic regime we find arrows of time with low intricacy giving rise to a time-local description of the approach to equilibrium as only a low finite number of derivatives of the autocorrelation function $C(t)$ at times 00 and $t$ are sufficient to construct an AOTF.
@@
in which we set $(J,\Delta^{\prime})=(1,0.5)$ and where $s_{k}^{a}=\sigma_{k}^{a}/2$ denote the spin-matrices. We vary the nearest-neighbour interaction term $\Delta$. Introducing the next-nearest interaction $\Delta^{\prime}$ breaks Bethe-Ansatz integrability [35, 36]. However, for $\Delta\gg\Delta^{\prime}$ the system again approaches the integrable regime, see Fig. 5. For any isotropies $\Delta,\Delta^{\prime}$ the total magnetization is conserved [35], motivating the study of the respective current
@@
Figure 4: Autocorrelation functions $C(t)$ (left) and their respective squares $C^{2}(t)$ and AOTFs $\alpha^{\mathcal{I}}$ (right) for the magnetisation current in the XXZ model (9) for different nearest-neighbour interactions $\Delta$. Top: At $\Delta=3$ (chaotic regime, see Fig. 5) $C(t)$ does not decay monotonously but the AOTF $\alpha^{3}(t)$ does. Bottom: Further away from the chaotic regime, at $\Delta=10$, an AOTF is reached only at $\mathcal{I}=47$, i.e. $\alpha^{47}(t)$ decays monotonously.
@@
Lastly, we include in our investigation the energy current
@@
where we choose the parameters as $(\omega,J,B_{x},B_{z})=(1,1,1,0.5)$. Varying the coupling strength $\lambda$ we compute the autocorrelation for the $x$ and $z-$component of coupled spin respectively, i.e. $\mathcal{O}^{(x)}=\sigma_{S}^{x}$ and $\mathcal{O}^{(z)}=\sigma_{S}^{z}$. See Fig. 8 for exemplary AOTFs of $\sigma_{S}^{z}.$
Figure 8: Autocorrelation functions $C(t)$ (left) and their respective squares $C^{2}(t)$ and AOTFs $\alpha^{\mathcal{I}}$ (right) for the observable $\sigma_{S}^{z}$ in the spin bath model (16) for different couplings $\lambda$. Top: At small coupling $\lambda$, here: $\lambda=2$ (chaotic regime, see Fig. 9) we find an AOTF with intricacy $\mathcal{I}=10$. Bottom: Approaching the chaotic regime again at $\lambda=12$, the intricacy of the respective AOTF drops to $\mathcal{I}=48$ again falling within our numerical scope.
@@
At small coupling $\lambda$ the intricacies $\mathcal{I}$ for the AOTFs of $\sigma_{S}^{z}$ start off at $\mathcal{I}=0$ as at first the autocorrelation functions are monotonous themselves. With increasing coupling strength, the intricacies grow until $\mathcal{I}=19$ for $\lambda=2.5$. As the system approaches the nonchaotic regime the intricacies exceed $\mathcal{I}=100$, leaving the scope of our numerical method. Once the system approaches the chaotic regime again at $\lambda\sim 11$ we again infer intricacies $\mathcal{I}<100$ starting at $\lambda=12$ where $\mathcal{I}=48$, cf. Fig. 8. These then decrease further as the system moves away from the nonchaotic regime, eventually giving rise to an AOTF with $\mathcal{I}=5$ at $\lambda=20$.
In the case of the observable $\sigma_{S}^{x}$ the picture is different. The intricacies do not seem to be affected by the systems’ transition from the chaotic to an nonchaotic regime. With an exception at $\lambda=2$ where our construction finds the respective autocorrelation function to be suitably monotonous, i.e. $\mathcal{I}=0$, we find AOTFs with an intricacy $\mathcal{I}=2$ for any other considered coupling strength $\lambda\leq 8.5$. Entering a strong coupling range at $\lambda\gtrsim 9$ however, the functions $C^{2}(t)$ no longer decay within a reasonable period in the sense we require to infer an equilibration time $\tau$. Consequently we leave them out of further discussions.
Having presented all numerical evidence, we now aim at distilling the common principle that is present in all these examples: Whenever an autocorrelation function of a few-body observable exhibits high intricacy, e.g. $\mathcal{I}>20$ (for the $\delta$ chosen here) the corresponding system either belongs to the nonchaotic regime (according to the gap ratio) or a nonchaotic regime is close by with respect to variation of some model parameter. Note that the converse does not hold: low-intricacy dynamics may emerge in nonchaotic regimes, see e.g. $\sigma_{S}^{x}$ in the spin-bath model. This principle is our main numerical result.